OpenStudy (anonymous):

a skateboard manufacturer finds that the cost c of making x skateboards per day is given by c(x) = x^2 -24x + 244 a) how many skateboards should be made per day to minimize the cost of production? b) what is the minimum cost? c) what is the cost if no skateboards were made in a day?

OpenStudy (anonymous):

well for part c it is -24 +244 so 244-24 (if that makes any sense??)

OpenStudy (unklerhaukus):

\[c(x)=x^2 -24x + 244 \] for part c) \[c(0)=(0)^2 -24(0) + 244 \]